Thanks In Advance!! :)
The frequency table shows the number of minutes that members of a fitness club spent exercising one weekend.

How many members spent fewer than 3 hours exercising?
Number of Minutes
Frequency
0–59 7
60–119 12
120–179 10
180–239 6
240–299 3
Answer Choices:
9
10 (I Think It's This One)
19
29

Answer by jim_thompson5910(35256)

3 hours = 3*60 = 180 minutes


Add up the frequencies that correspond to minutes that span from 0 to 180 (don't include 180): 7+12+10 = 19+10 = 29

So the answer is actually 29